Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0. Solution:10,only solution.
Answer by ankor@dixie-net.com(22740)

Write a quadratic equation in the variable x having the given numbers as solutions.
Type the equation in standard form, ax^2+bx+c=0. Solution:10,only solution.
x=10
x=10
It will be a perfect square with factors
(x-10)(x-10) = 0
FOIL
f(x) = x^2 -10x - 10x + 100
combine like terms
f(x) = x^2 - 20x + 100